OpenStudy (anonymous):

I need to to show that, using the quotient rule: g(x)= (e(-3x+2sinx)/6) ---------------- (3+2cosx) has derivative g'(x)= - (4(sin^2)x -12sin x +5) (e(-3x+2sinx)/6) ----------------------------------- 6(3+2cosx)^2

OpenStudy (mertsj):

\[g(x)=\frac{e ^{\frac{-3x+2\sin x}{6}}}{3+2\cos x}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{u}{v}=\frac{u'v-v'u}{v^2}\]

OpenStudy (anonymous):

\[(\frac{u}{v})'\]

OpenStudy (anonymous):

The problem is to simplify to get the deritive as expressed. Could you show me the complete working out

OpenStudy (anonymous):

Denominator = \[(3+2\cos x)^2\] Numerator : \[u'=(\frac{-1}{2}+\frac{1}{3}\cos x)e^{\frac{-3x+2\sin x}{6}}\] \[v'=-2\sin x\] Notice that : \[(\frac{-1}{2}+\frac{1}{3}\cos x)(3+2\cos x)=\frac{4\cos^2 x - 9}{6}\] Factor out \[e^{\frac{-3x+2\sin x}{6}}\] Thus numerator = \[e^{\frac{-3x+2\sin x}{6}}(\frac{4\cos^2 x-9}{6}+2\sin x)\] Notice again : \[\cos^2 x +\sin^2 x=1\] Hence numerator = \[e^{\frac{-3x+2\sin x}{6}}(\frac{-5-4\sin^2 x+12\sin x}{6})\] .... :)

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (anonymous):

you're welcome ;)

OpenStudy (anonymous):

Hi anhkoavo1210

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Ok i'm here

OpenStudy (anonymous):

which part in numerator you want I explain?

OpenStudy (anonymous):

Could you show me how the how I get from e(-3x+2sinx)/6) to - (4(sin^2)x -12sin x +5) (e(-3x+2sinx)/6)

OpenStudy (anonymous):

I can do some of it with the quotient rule but get stuff in the simpplification process

OpenStudy (anonymous):

get stuck with the simplification process

OpenStudy (anonymous):

uh, can you display your numerator you're having

OpenStudy (anonymous):

I get to this point (2cosx+3)(e^(2sinx-3x)/6)(2cosx-3) - (e^(2sinx-3x))(-2sinx) --------------------- 6 Then i get stuck

OpenStudy (anonymous):

2cosx+3)(e^(2sinx-3x)/6)(2cosx-3) - (e^(2sinx-3x)/6)(-2sinx) --------------------- 6

OpenStudy (anonymous):

that is what it should be - sorry typing is difficult

OpenStudy (anonymous):

Ok, I write down again : \[(\frac{-1}{2}+\frac{1}{3} \cos x)e^{\frac{−3x+2sinx}{6}}(3+2\cos x)+e^{\frac{−3x+2sinx}{6}}2\sin x\] You got it here?

OpenStudy (anonymous):

How did you get -1/2 + 1/3 cosx

OpenStudy (anonymous):

You try to see \[(e^{\frac{−3x+2sinx}{6}})'=?\]

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

And how to I get to my derivative using cos^2+sin^2x = 1

OpenStudy (anonymous):

for the last part as you mentioned before

OpenStudy (anonymous):

Did you got it here? \[e^{\frac{−3x+2sinx}{6}}(\frac{4\cos^2 x−9}{6}+2\sin x)\]

OpenStudy (anonymous):

Can explain how you do this part?

OpenStudy (anonymous):

you see\[ (−\frac{1}{2}+\frac{1}{3}\cos x)(3+2cosx)=\frac{4\cos^2x−9}{6}\] ??

OpenStudy (anonymous):

you factored?

OpenStudy (anonymous):

expanded

OpenStudy (anonymous):

no, you see \[\frac{−1}{2}+\frac{1}{3}\cos x=\frac{-3+2\cos x}{6}\] After I use formula : \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Thanks for you help. I have to sit down and spend time on this. Thanks very much. I am making some progress. Do you min if I as for some help in the future. I am also and English teacher so if you need help from me then I more then happy to help, even conversation where I can correct you English expression even over chat. Let me know. I have to go now.

OpenStudy (anonymous):

I meant if you mind if I ask for more help in the next few days as I have an exam soon.

OpenStudy (anonymous):

:D ok, thank you. Do you have a nickname yahoo?

OpenStudy (anonymous):

Are you on skype?

OpenStudy (anonymous):

No i usually use yahoo.

OpenStudy (anonymous):

this is my email address on yahoo hussain_m_ismail@yahoo.co.uk

OpenStudy (anonymous):

tell me your name :D

OpenStudy (anonymous):

I have to go now so keep in touch. And I will contact you if i need more help, if that is ok

OpenStudy (anonymous):

Ok, if I available. But your name is?